What is the effective resistance for a parallel circuit containing a 12 V potential and three resistors R1= 3 ohms,

R2=2 ohms and R3= 1 ohm.


1
Expert's answer
2021-05-10T09:22:57-0400

Gives

R1=3ohm

R2=2ohm

R3=1ohm


Potential (V)=12 V

Rnet=R1||R2||R3

"\frac{1}{R_{net}}=\frac{11}{6}"

Rnet="\frac{6}{11}ohm"

"I=\frac{V}{R}=\frac{12}{\frac{6}{11}}=22Amp"